# Pseudo-Autoregressive Neural Codec Language Models for Efficient Zero-Shot Text-to-Speech Synthesis

### v1

@@ 5.1. Experimental Setup
Model Configurations Both stages of PALLE adopt the same model architecture, comprising: a decoder-only Transformer with 12 layers, 16 attention heads, 1024-dimensional embeddings, and 4096-dimensional feed-forward layers; a one-layer ConvNeXt V2 block with 1024-dimensional embeddings and a 2048-dimensional feed-forward layer; and a convolutional positional embedding with a kernel size of 7; totaling 177.0M parameters. For text tokenization, a 2,000-class BPE is used. Speech tokenization is carried out using the off-the-shelf S3Tokenizer [^2] https://github.com/xingchensong/S3Tokenizer from CosyVoice 2 (Du et al., 2024c) at a token rate of 25 Hz. Speech reconstruction is performed using the off-the-shelf CFM model with a built-in vocoder, also from CosyVoice 2 (Du et al., 2024c).
@@
Inference Details The first stage runs 100 steps per sample of length $T$, each generating a span of size $k=\min(\left\lfloor 0.01T\right\rfloor,N_{\text{left}})$, which corresponds to a fixed percentage of the sample length and is smaller than the span size used during training [^3] Although the model is trained with span size of $\left\lfloor 0.1T\right\rfloor$, we found the best configuration for inference is $\left\lfloor 0.01T\right\rfloor$, as detailed in Sec.  5.6.. A detailed analysis of the per-step prediction ratio is provided in Fig. 4 of the ablation study. The second stage runs 7 steps per sample using the $\gamma=0.05$ quantile of the confidence scores. Top-$p$ sampling with a small $p$ in the range of 0.2 to 0.35 yields slight performance improvements over greedy search, without introducing much randomness or additional computational cost.
@@ 5.2. Evaluation
Evaluation Settings We use the LibriSpeech (Panayotov et al., 2015) test-clean set for zero-shot TTS evaluation, ensuring no speaker overlap with training data. Following previous works (Wang et al., 2023; Chen et al., 2025), we use the same test set, which consists of segments ranging from 4 to 10 seconds, totaling 2.2 hours of data from 40 unique speakers and 1,234 samples. Following previous works (Wang et al., 2023; Chen et al., 2025), our evaluations contain two inference tasks:
@@
  • Cross-Sentence: Given a reference utterance and its transcription as the prompt, the model synthesizes speech for the target text while preserving the speaker’s characteristics.
@@
  • WER (Word Error Rate) measures the robustness and intelligibility of synthesized speech. Neural TTS systems sometimes suffer from robustness issues, including word deletion, insertion, substitution, or generating endless silence or noise, due to attention misalignments. To evaluate these aspects, ASR models are used to transcribe the synthesized audio. WER is computed between the generated transcripts and the ground-truth text, denoted as WER-W and WER-H, with Whisper Large-v3444https://huggingface.co/openai/whisper-large-v3 (Radford et al., 2023) and the HuBERT-Large ASR model [^5] https://huggingface.co/facebook/hubert-large-ls960-ft (Hsu et al., 2021) respectively.
@@
  • VALL-E (Wang et al., 2023): A two-stage TTS system that utilizes both AR and NAR models to predict RVQ tokens at 75Hz from EnCodec (Défossez et al., 2023). We evaluate three versions of VALL-E: (1) the model trained on Librilight (Kahn et al., 2020), and (2) the model trained on Libriheavy (Kang et al., 2024), using performance metrics reported in (Wang et al., 2023; Chen et al., 2025) and RTF taken from (Meng et al., 2024) measured using the official checkpoint; and (3) the model trained on LibriTTS (Zen et al., 2019), which we reproduce using the open-source implementation [^7] https://github.com/lifeiteng/vall-e.
  • CosyVoice (Du et al., 2024a): A two-stage TTS system that utilizes a text-to-token AR model and a token-to-speech flow-matching model. We compare against the 25Hz version trained on LibriTTS using the official code, with results reported in (Sheng et al., 2025).
  • CosyVoice 2 (Du et al., 2024c): A two-stage TTS system that utilizes an interleaved text-speech AR model and a chunk-aware token-to-speech flow-matching model. We compare against the 25Hz version trained on LibriTTS, with results reported in (Sheng et al., 2025).
  • MaskGCT (Wang et al., 2024): A fully NAR two-stage TTS system based on mask-and-predict learning, consisting of a 695M text-to-semantic model and a 353M semantic-to-acoustic model. We use the official pre-trained checkpoint [^8] https://huggingface.co/amphion/MaskGCT, trained on 100K hours of in-the-wild Chinese and English speech data from the Emilia (He et al., 2024) dataset.
  • E2 TTS (Eskimez et al., 2024): A fully NAR TTS system based on flow matching, with a model size of  [^3] M. We use the open-source pre-trained checkpoint [^9] https://huggingface.co/SWivid/E2-TTS, trained on 100K hours of in-the-wild Chinese and English speech data from Emilia (He et al., 2024).
  • F5-TTS (Chen et al., 2024b): A fully NAR TTS system based on flow matching, with a model size of 336M. We use the official pre-trained checkpoint [^10] https://huggingface.co/SWivid/F5-TTS, trained on 100K hours of in-the-wild Chinese and English speech data from Emilia (He et al., 2024).
@@ 5.3. Evaluation Results
5.5. Total Duration Impact Analysis
We evaluate the robustness of the stage-one and two-stage variants of PALLE under varying total duration multipliers on the cross-sentence task, where each multiplier corresponds to a different speech tempo. As shown in Fig. 5, the duration of the generated speech is scaled from 0.7$\times$ to 1.3$\times$ of the ground-truth duration. Both systems achieve the lowest WER-H at a multiplier of 1.1 and the highest SIM at 1.2, indicating that generating speech at a slightly slower pace than the ground truth improves both intelligibility and speaker similarity. One possible explanation is that, in the cross-sentence task, the speaking rate of the speech prompt may differ from that of the target. When the prompt is substantially faster, the model may struggle to maintain prosodic consistency via in-context learning, leading to degraded synthesis quality. Overall, both systems exhibit stable performance across a moderate tempo range, with near-optimal results between 0.9$\times$ and 1.3$\times$, indicating robustness to speaking rate variation.
@@ 5.6. Ablation Study
Joint Two-Stage Modeling We attempt to unify stage one and stage two into a single model via multitask finetuning, randomly performing either task per sample with equal probability. While performance slightly decreases on the continual task, WER on the cross-sentence task increases by 20%, regardless of whether ground-truth durations are used. This may be attributed to the halved parameter capacity per stage compared to the separate models.
@@ 6. Conclusions
In this paper, we propose a novel sequence generation paradigm, pseudo-autoregressive language modeling (PAR), which combines the strengths of temporal modeling from AR and parallel decoding from NAR, and generates dynamic-length spans at fixed time steps. We apply PAR to zero-shot TTS tasks, resulting in a new TTS model called PALLE. PALLE leverages a PAR model to generate initial discrete speech tokens with fixed steps for various target utterances, followed by refinement using a NAR model. Experimental results demonstrate that PALLE significantly improves intelligibility, speaker similarity, and efficiency. Looking ahead, PAR holds potential for application in broader generation domains, such as text, image, and video.

### v2

@@ 5.1. Experimental Setup
Model Configurations Both stages of PALLE adopt the same model architecture, comprising: a decoder-only Transformer with 12 layers, 16 attention heads, 1024-dimensional embeddings, and 4096-dimensional feed-forward layers; a one-layer ConvNeXt V2 block with 1024-dimensional embeddings and a 2048-dimensional feed-forward layer; and a convolutional positional embedding with a kernel size of 7; totaling 177.0M parameters. For text tokenization, a 2,000-class BPE is used. Speech tokenization is carried out using the off-the-shelf S3Tokenizer [^2] https://github.com/xingchensong/S3Tokenizer from CosyVoice 2 (Du et al., 2024b) at a token rate of 25 Hz. Speech reconstruction is performed using the off-the-shelf CFM model with a built-in vocoder, also from CosyVoice 2 (Du et al., 2024b).
@@
Inference Details The first stage runs 100 steps per sample of length $T$, each generating a span of size $k=\min(\left\lfloor 0.01T\right\rfloor,N_{\text{left}})$, which corresponds to a fixed percentage of the sample length and is smaller than the span size used during training [^3] Although the model is trained with span size of $\left\lfloor 0.1T\right\rfloor$, we found the best configuration for inference is $\left\lfloor 0.01T\right\rfloor$, as detailed in Sec.  5.5.. A detailed analysis of the per-step prediction ratio is provided in Fig. 4 of the ablation study. The second stage runs 7 steps per sample using the $\gamma=0.05$ quantile of the confidence scores. Top-$p$ sampling with a small $p$ in the range of 0.2 to 0.35 yields slight performance improvements over greedy search, without introducing much randomness or additional computational cost.
@@ 5.2. Evaluation
Evaluation Settings We use the LibriSpeech (Panayotov et al., 2015) test-clean set for zero-shot TTS evaluation, ensuring no speaker overlap with training data. Following previous works (Wang et al., 2023; Chen et al., 2025b), we use the same test set, which consists of segments ranging from 4 to 10 seconds, totaling 2.2 hours of data from 40 unique speakers and 1,234 samples. Our evaluations contain two inference tasks:
@@
  • Cross-Sentence: Given a reference utterance and its transcription as the prompt, the model synthesizes speech for the target text.
@@
  • WER (Word Error Rate) measures the robustness and intelligibility of synthesized speech. WER is computed between ASR transcription of synthesized audio and the ground-truth text, denoted as WER-W and WER-H, with Whisper Large-v3444https://huggingface.co/openai/whisper-large-v3 (Radford et al., 2023) and the HuBERT-Large ASR model [^5] https://huggingface.co/facebook/hubert-large-ls960-ft (Hsu et al., 2021) respectively.
@@
  • VALL-E (Wang et al., 2023): A two-stage AR+NAR system predicting 75Hz EnCodec RVQ tokens (Défossez et al., 2023). We evaluate three variants trained on Librilight (Kahn et al., 2020), Libriheavy (Kang et al., 2024) and LibriTTS (Zen et al., 2019), with RTF from (Meng et al., 2025) using the official checkpoint.
  • CosyVoice (Du et al., 2024a): A two-stage AR+flow-matching system. We evaluate the 25Hz LibriTTS version, with results from (Sheng et al., 2025).
  • CosyVoice 2 (Du et al., 2024b): A two-stage interleaved AR+chunk-aware flow-matching system. We evaluate the 25Hz LibriTTS version, with results from (Sheng et al., 2025).
  • MaskGCT (Wang et al., 2024): A two-stage NAR mask-and-predict system with 695M text-to-semantic and 353M semantic-to-acoustic models. We evaluate the official checkpoint [^7] https://huggingface.co/amphion/MaskGCT trained on 100K hours of Chinese and English speech from Emilia (He et al., 2024).
  • E2 TTS (Eskimez et al., 2024) & F5-TTS (Chen et al., 2025a): Fully NAR flow-matching systems with  [^3] M and 336M parameters. We evaluate open-source checkpoints [^8] https://huggingface.co/SWivid/E2-TTS [^9] https://huggingface.co/SWivid/F5-TTS trained on 100K hours of bilingual Emilia (He et al., 2024).
@@ 5.5. Ablation Study
Total Duration We evaluate the robustness of the stage-one and two-stage variants of PALLE under varying total duration multipliers on the cross-sentence task, where each multiplier corresponds to a different speech tempo. As shown in Fig. 5, the duration of the generated speech is scaled from 0.7$\times$ to 1.3$\times$ of the ground-truth duration. Both systems achieve the lowest WER-H at a multiplier of 1.1 and the highest SIM at 1.2, indicating that generating speech at a slightly slower pace than the ground truth improves both intelligibility and speaker similarity. One possible explanation is that, in the cross-sentence task, the speaking rate of the speech prompt may differ from that of the target. When the prompt is substantially faster, the model may struggle to maintain prosodic consistency via in-context learning, leading to degraded synthesis quality. Overall, both systems exhibit stable performance across a moderate tempo range, with near-optimal results between 0.9$\times$ and 1.3$\times$, indicating robustness to speaking rate variation.
@@
Joint Two-Stage Modeling We attempt to unify stage one and stage two into a single model via multitask finetuning, randomly performing one task per sample. However, stage two loss degrades stage one performance. While performance slightly decreases on the continuation task, WER on the cross-sentence task increases by 20%, regardless of whether ground-truth durations are used. This may be attributed to the halved parameter capacity per stage compared to the separate models.
@@ 6. Conclusions
In this paper, we propose a novel sequence generation paradigm, pseudo-autoregressive language modeling (PAR), which combines the strengths of temporal modeling from AR and parallel decoding from NAR, and generates dynamic-length spans at fixed time steps. We apply PAR to zero-shot TTS tasks, resulting in a new TTS model called PALLE. PALLE leverages a PAR model to generate initial discrete speech tokens with fixed steps for various target utterances, followed by refinement using a NAR model. Experimental results demonstrate that PALLE significantly improves intelligibility, speaker similarity, and efficiency. Looking ahead, PAR holds potential for broader applicability in generation domains, such as text, image, and video.
7. Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. 62206171 and No. U23B2018), Shanghai Municipal Science and Technology Major Project under Grant 2021SHZDZX0102 and Yangtze River Delta Science and Technology Innovation Community Joint Research Project (2024CSJGG01100).